1. The dimerization of ClO is second order in ClO.
2ClO(g) ⇌ Cl2O2
Time (s) [ClO] (molecules/cm3)
0 2.08 x 105
1 8.64 x 104
2 5.46 x 104
3 3.99 x 104
a) Use the following data to determine the value of k at 350 K. answer in cm3/molecules•s
b) Use the data above to determine the value the half-life at 350 K. answer in s
1
Expert's answer
2020-03-27T09:00:21-0400

The rate constant for the given reaction is calculated as shown below.

a) If the reaction refers to second-order reactions (С01 = С02), then the rate constant:


k=1/t(1/c-1/c0)

k=1/1(1/8,64 х 104-1/2.08 x 105) =1(1/8,64 х 104-1/20.8 x 104)=(20.8x 104-8,64 х 104)/20.8x 104х8,64 х 104=12.16х 104/179.712х 104=0.0676х 104 cm3/molecules•s 

k=1/2(1/5.46 x 104-1/ 2.08 x 105)=1/2(20.8x 104-5,46х 104)/20.8x 104х4,46х104)=1/215.34х 104/113.568х 104=1/2 х 0.1350х 104/molecules•s=0.0676 х 104/molecules•s

k=1/3(1/3.99 x 104-1/ 2.08 x 105)=1/3(20.8x 104-3.99х 104)/20.8x 104х3.99х104)=1/316.81х 104/82.992х 104 =1/3 0.2025 х 104/molecules•s=0.0676х 104 cm3/molecules•s


b) t1/2=1/k х 1/c0=1/0.0676х 104 cm3/molecules•sх 1/20.8 x 104cm3/molecules•s)=1/1.40608x104=0.71x10-4s